# Is Seeing Believing? Evaluating Human Sensitivity to Synthetic Video

### 14 Mar 2026 v1

@@ 1 Introduction
Recent advances in machine learning architectures capable of synthesizing highly realistic video and audio material, so-called deepfakes, threaten the integrity of videos as reliable information sources Westerlund (2019). A central concern is that malicious actors can use the technology to create and spread false evidence, as deepfakes enable the creation of convincing impersonations of influential and trusted figures Köbis et al. (2021); Mink et al. (2024), and are frequently deployed in contexts of sexual harassment Wolfe et al. (2023); Gebru and Denton (2024). The introduction of deepfake videos raises concerns as people seem to be particularly susceptible to video-based disinformation Sundar et al. (2021); Lee and Shin (2022). Despite the growing awareness of issues concerning deepfakes, we lack a general understanding of how people perceive and evaluate synthetically generated media Tahir et al. (2021); Gamage et al. (2022).
@@
The visual aspect of a video may influence how its message is perceived in several ways: It may contain arguments, focus or distract the audience’s attention, and serve as cue for heuristic evaluations Peng et al. (2023). A key heuristic in processing visual information is processing fluency, the subjective ease of processing a stimulus. When information is easier (or harder) to process than expected, this metacognitive “feel” systematically shifts judgments such as truth, trust, and liking Whittlesea and Williams (2000). Prior research has found evidence for the influence of processing fluency on judgments of credibility and trust in numerous media types Schwarz et al. (2021). The legibility of text Reber and Schwarz (1999), the quality of audio recordings Newman and Schwarz (2018), and the presence of nonprobative photos alongside a statement Newman et al. (2012) were found to affect how credible people deemed information to be. Nevertheless, the effects of processing fluency of videos remain mostly underexplored.
@@
Given the unique characteristics of videos and the increasing prevalence of deepfakes, studying the effects of visual and auditory video aspects on credibility becomes crucial Gamage et al. (2022). This is because artifacts remain the major indicator of artificially generated content, despite the increasing quality of deepfake videos Ciftci et al. (2020). These artifacts often manifest as unnatural movements, extraneous limbs, inconsistencies in the subject’s lighting, and similar features that do not occur in recorded videos Matern et al. (2019).
@@
Distortions and deepfake artifacts can reduce processing fluency because they violate viewers’ expectations about how recorded videos typically look and sound. Such low-level irregularities act as perceptual expectancy violations that make the video subjectively harder to process, and may therefore serve as heuristic cues in credibility judgments. This is especially relevant in online video environments, where high information density and limited attention promote intuitive, heuristic modes of evaluation Greifeneder et al. (2011); Lang (2000). Under these conditions, which are common on platforms such as YouTube and TikTok Metzger et al. (2010); Sundar et al. (2021), viewers are likely to rely on directly accessible perceptual cues rather than analytical assessment. We therefore examine whether visual distortions, auditory distortions, and deepfake artifacts in faces and voices systematically affect video credibility, processing fluency, and related evaluation outcomes.
@@
For HCI, credibility and authenticity are interactional judgments shaped by interface cues as much as by content. Online videos are typically encountered in environments that impose cognitive load and encourage heuristic processing Lang (2000); Metzger and Flanagin (2015), meaning that subtle audiovisual irregularities become salient inputs to users’ credibility assessments. Yet, we lack empirical evidence on how users in realistic viewing conditions respond to artifacts in synthetic media Hilligoss and Rieh (2008); Sundar (2008); Metzger and Flanagin (2015).
@@ 2.1 Deepfakes and Artifacts
Deepfakes are videos in which faces and/or voices have been synthetically generated using deep-learning models. The technology enables convincing impersonations with comparatively little expertise Kaate et al. (2023); Roose (2018) and is increasingly used in deceptive contexts Dobber et al. (2021). For the present work we distinguish three categories: (1) recorded videos, captured without AI synthesis; (2) distorted videos, where visual or auditory degradations are added without using generative models; and (3) deepfake videos, in which the facial appearance and/or narration have been synthesized.
Deepfake generation pipelines introduce characteristic artifacts because models must estimate facial geometry, motion, illumination, and texture, and temporally align synthesized audio and video. Imperfections in these steps produce residual cues that differentiate synthetic from recorded footage, such as subtle inconsistencies in shading, boundary alignment, facial dynamics, or mouth–speech synchrony Matern et al. (2019); Li et al. (2020); Yang et al. (2019); Ciftci et al. (2020). While some artifacts may be imperceptible to viewers but detectable algorithmically Agarwal et al. (2020), others manifest as visible irregularities or atypical movement patterns.
Despite these cues, viewers often struggle to reliably distinguish deepfakes from recorded videos. Prior studies show that many people fail to identify deepfakes even when aware of their existence Korshunov and Marcel (2021); Groh et al. (2022); Köbis et al. (2021); Mink et al. (2024), and deepfake political videos frequently receive moderate credibility ratings Lee and Shin (2022); Hwang et al. (2021). Uncertainty is widespread, and judgments are easily confounded by factors such as familiarity, liking, or sympathy for the depicted individual Vaccari and Chadwick (2020); Köbis et al. (2021). This difficulty underscores the need to understand how specific kinds of visual and auditory artifacts shape users’ evaluations.
@@ 2.2 Heuristics and processing fluency in credibility judgments
Credibility refers to an individual’s judgment of the veracity of the content of communication Appelman and Sundar (2016). Online environments complicate such judgments: information volume is high, source cues are sparse or ambiguous, and professional verification is limited Callister (2000); Metzger (2007); Metzger and Flanagin (2015); Sohrawardi et al. (2024). Under these conditions of overload and limited cognitive resources, people rely on heuristics; that is, readily accessible cues that can be evaluated quickly but are fallible Lang (2000); Hilligoss and Rieh (2008). Social indicators such as endorsements, comments, or source reputation often serve this function Metzger et al. (2010), and popularity has been shown to be influential for credibility assessments explicitly in the context of deepfake content Jin et al. (2025).
Deepfakes complicate credibility judgments further because they can appropriate trusted identities and obscure more diagnostic cues. Although synthetic videos may appear plausible, they often contain subtle irregularities. Such unexpected visual or auditory features create expectancy violations; deviations from what viewers assume recorded videos should look or sound like, which can reduce processing fluency, the subjective ease of processing a stimulus Whittlesea and Williams (2000). Processing fluency reliably influences credibility judgments across media: less legible text Reber and Schwarz (1999) and lower-quality audio Newman and Schwarz (2018) reduce perceived truthfulness Schwarz et al. (2021).
People particularly rely on fluency when external knowledge is limited Alter and Oppenheimer (2009). One proposed mechanism is familiarity: fluently processed information feels more familiar and is therefore judged as more likely to be true Dechêne et al. (2010); Whittlesea et al. (1990). In the context of deepfakes, this suggests that viewers may turn to the “feel” of ease or difficulty when evaluating unfamiliar videos or presenters. However, it remains unclear whether and how such fluency-based heuristics operate for videos, and especially for deepfake-generated faces and voices, where multiple perceptual channels may produce competing or ambiguous cues.
@@ 2.3 Credibility assessment as an interaction problem
In online environments, credibility judgments are shaped by an interactional process that includes the viewer and the design of digital interfaces. Users evaluate authenticity within feeds, video players, and recommendation systems that structure attention and impose cognitive constraints Hilligoss and Rieh (2008); Metzger and Flanagin (2015). HCI research has shown that interface-level cues such as modality, visual clarity, and system agency systematically influence heuristic credibility judgments Sundar (2008); Sun et al. (2024); Zieglmeier and Lehene (2021). In this perspective, deepfake artifacts can be understood as perceptual cues embedded in the interaction environment: they are surfaced to the user through the interface in the same way as other credibility-relevant signals. Studying how users respond to these artifacts therefore addresses the central question of how interactive systems mediate trust and uncertainty when synthetic media are encountered in real-world interface contexts.
@@ 4.4 Dependent variables
Following Graf et al. (2018), subjective processing fluency was measured with a slider, with internally stored responses between 0 (difficult) and 100 (easy). This measure was used in all three studies.
@@
Based on the scale by Appelman and Sundar (2016), message credibility was computed as the mean of three 7-point Likert items accurate, authentic, and believable, with response options ranging from 1 (“describes very poorly”) to 7 (“describes very well”). This scale was used in all three studies.
@@
In Study II and Study III, conceptual and factual learning were measured as the pre-to-post difference in a multiple choice questionnaire, following guidelines by Petersen et al. (2021). Pre- and post-stimulus knowledge questionnaires were administered with mixed item order and embedded attention checks.
@@ 4.5 Stimuli
An educational video was produced to serve as a common stimulus source for all three studies. The video was $447\text{\,}\mathrm{s}\mathrm{e}\mathrm{c}\mathrm{o}\mathrm{n}\mathrm{d}\mathrm{s}$ long and featured an actor presenting a script adapted from Petersen et al. (2021). All video editing was performed in Adobe After Effects 23.2.1 (see Appendix 2 and Appendix 3 for details on video processing). The original script consisted of sections on general virology, measles, Zika, and COVID-19; we excluded the COVID-19 section to avoid interference from prevalent, non-scientific interpretations of the disease.
@@
We used the Latent Image Animator with “vox-pt” weights by Wang and colleagues Wang et al. (2022) to synthesize a deepfake version of the video. This approach requires a $256\times 256$ close-up video of a face and a $256\times 256$ portrait picture to generate a $256\times 256$ video of the portrait picture exhibiting the behavior captured in the video. The resulting animated portrait was then pasted into and aligned with the source video to obtain the deepfaked stimulus video. We synthesized a version of the presenter’s narration by fine-tuning a So-Vits-Svc-fork model [^2] https://github.com/voicepaw/so-vits-svc-fork. The model was trained on a $14.25\text{\,}\mathrm{m}\mathrm{i}\mathrm{n}\mathrm{u}\mathrm{t}\mathrm{e}\mathrm{s}$ long recording of a diverse set of English sentences adapted from the Harvard sentences [^3] https://harvardsentences.com. Once trained and provided with the audio of the source video, the model synthesized a version of the original narration in the new voice. The synthesized video stimulus was created by replacing the presenter’s face in the source video with the deepfake animation while retaining the original audio. The synthesized narration stimulus replaced the source video’s original narration track with the synthetic version. For the synthesized video and narration condition, we combined the deepfake stimulus video with the synthesized narration track.
@@ 5.1 Study I
More participants detected a digital alteration in the reduced visual clarity condition ($n=126$) than in the baseline condition ($n=22$). More participants were uncertain about digital alterations in the baseline group ($n=59$) than in the reduced visual clarity group ($n=32$). Similarly, more participants in the baseline group did not detect an alteration ($n=81$) compared to the reduced visual clarity group ($n=13$; see Figure 4). Building on work by Altay and Gilardi (2024), who found that the belief of text being AI-generated negatively impacting its perceived credibility, we expanded on our preregistered analysis and examined whether believing that the video was digitally altered was associated with lower credibility. A Spearman rank correlation indicated a moderate negative association between alteration detection and credibility (r(331)=.25,p<.001r(331)=-.25,\,p<.001), see Appendix 7 for details.
@@ 5.3 Study III
Figure 7: Distribution of credible differences between the baseline video and the fully synthesized video. (a) shows the distribution of message credibility differences, (b) the distribution for processing fluency in Study III. Bayesian estimation Kruschke (2013) was used to calculate distributions of mean processing fluency and message credibility values for the baseline and synthesized video and narration conditions. The differences between the means are expressed in fractions of the pooled standard deviation (i.e., in units comparable to Cohen’s $d$). The 95% HDIs (High Density Intervals) indicate where the bulk of the credible differences fall. For processing fluency (b), the HDI is close to or includes 0, consistent with no meaningful difference between conditions. For message credibility (a), the HDI lies substantially above 0, and essentially all credible effect sizes exceed a small-effect threshold (Cohen’s $d=0.10$), indicating that the distance between the means is too large to accept $H_{0}$.
@@ 6 Discussion
In Study II, visual distortions and audio-video asynchrony reduced credibility compared to audio distortions alone but not the baseline, highlighting the role of sensory cues in credibility assessment. Study III found that only the combination of deepfake visuals and narration significantly reduced credibility compared to the baseline, suggesting a small additive effect. Contrary to Newman and Schwarz Newman and Schwarz (2018), we found no significant impact of audio manipulations on credibility.
@@
Groups of participants that indicated having and not having detected an alteration were present across all three studies, regardless of the presence of an manipulation. The data reveals a significant rate of inaccuracy in identifying distortions and deepfake videos, reflected in the high proportion of participants selecting “I don’t know.” These findings are consistent with Vaccari and Chadwick’s research, which links deepfakes to increased uncertainty Vaccari and Chadwick (2020). Interestingly, the baseline condition in our studies highlights that the absence of manipulations does not reduce uncertainty. In Studies II and III, uncertainty was lowest in manipulation conditions and highest in baseline conditions without manipulations. Across studies, detecting (or believing to detect) an alteration was consistently associated with lower credibility. The negative correlations were small to moderate in Studies I and II and largest in Study III, indicating that perceived alteration reduced credibility regardless of whether manipulations were actually present.
@@
Our findings are in line with the association between source vividness and credibility as discussed by Lee and Shin in the context of deepfakes and misinformation Lee and Shin (2022). We find, however, our source vividness measure to be more affected than our message credibility measure. This aligns with the understanding of video credibility assessment as complex and affected by multiple factors. While all participants learned from exposure to the stimulus, none of the employed manipulations affected their learning success. All learning-relevant information was mediated via the audio track of the videos. This suggests that the audio-based manipulations employed in Study II and Study III did not interfere with comprehensibility and that the measured reduction in source vividness and message credibility did not coincide with reduced apprehension.

### 17 Mar 2026 v2

@@ 1 Introduction
Recent advances in machine learning architectures capable of synthesizing highly realistic video and audio material, so-called deepfakes, threaten the integrity of videos as reliable information sources [46]. A central concern is that malicious actors can use the technology to create and spread false evidence, as deepfakes enable the creation of convincing impersonations of influential and trusted figures [19, 29], and are frequently deployed in contexts of sexual harassment [49, 11]. The introduction of deepfake videos raises concerns as people seem to be particularly susceptible to video-based disinformation [40, 23]. Despite the growing awareness of issues concerning deepfakes, we lack a general understanding of how people perceive and evaluate synthetically generated media [43, 10].
@@
The visual aspect of a video may influence how its message is perceived in several ways: It may contain arguments, focus or distract the audience’s attention, and serve as cue for heuristic evaluations [33]. A key heuristic in processing visual information is processing fluency, the subjective ease of processing a stimulus. When information is easier (or harder) to process than expected, this metacognitive “feel” systematically shifts judgments such as truth, trust, and liking [48]. Prior research has found evidence for the influence of processing fluency on judgments of credibility and trust in numerous media types [37]. The legibility of text [35], the quality of audio recordings [31], and the presence of nonprobative photos alongside a statement [30] were found to affect how credible people deemed information to be. Nevertheless, the effects of processing fluency of videos remain mostly underexplored.
@@
Given the unique characteristics of videos and the increasing prevalence of deepfakes, studying the effects of visual and auditory video aspects on credibility becomes crucial [10]. This is because artifacts remain the major indicator of artificially generated content, despite the increasing quality of deepfake videos [7]. These artifacts often manifest as unnatural movements, extraneous limbs, inconsistencies in the subject’s lighting, and similar features that do not occur in recorded videos [25].
@@
Distortions and deepfake artifacts can reduce processing fluency because they violate viewers’ expectations about how recorded videos typically look and sound. Such low-level irregularities act as perceptual expectancy violations that make the video subjectively harder to process, and may therefore serve as heuristic cues in credibility judgments. This is especially relevant in online video environments, where high information density and limited attention promote intuitive, heuristic modes of evaluation [13, 22]. Under these conditions, which are common on platforms such as YouTube and TikTok [26, 40], viewers are likely to rely on directly accessible perceptual cues rather than analytical assessment. We therefore examine whether visual distortions, auditory distortions, and deepfake artifacts in faces and voices systematically affect video credibility, processing fluency, and related evaluation outcomes.
@@
For HCI, credibility and authenticity are interactional judgments shaped by interface cues as much as by content. Online videos are typically encountered in environments that impose cognitive load and encourage heuristic processing [22, 27], meaning that subtle audiovisual irregularities become salient inputs to users’ credibility assessments. Yet, we lack empirical evidence on how users in realistic viewing conditions respond to artifacts in synthetic media [15, 41, 27].
@@ 2.1 Deepfakes and Artifacts
Deepfakes are videos in which faces and/or voices have been synthetically generated using deep-learning models. The technology enables convincing impersonations with comparatively little expertise [18, 36] and is increasingly used in deceptive contexts [9]. For the present work we distinguish three categories: (1) recorded videos, captured without AI synthesis; (2) distorted videos, where visual or auditory degradations are added without using generative models; and (3) deepfake videos, in which the facial appearance and/or narration have been synthesized.
Deepfake generation pipelines introduce characteristic artifacts because models must estimate facial geometry, motion, illumination, and texture, and temporally align synthesized audio and video. Imperfections in these steps produce residual cues that differentiate synthetic from recorded footage, such as subtle inconsistencies in shading, boundary alignment, facial dynamics, or mouth–speech synchrony [25, 24, 50, 7]. While some artifacts may be imperceptible to viewers but detectable algorithmically [1], others manifest as visible irregularities or atypical movement patterns.
Despite these cues, viewers often struggle to reliably distinguish deepfakes from recorded videos. Prior studies show that many people fail to identify deepfakes even when aware of their existence [20, 14, 19, 29], and deepfake political videos frequently receive moderate credibility ratings [23, 16]. Uncertainty is widespread, and judgments are easily confounded by factors such as familiarity, liking, or sympathy for the depicted individual [44, 19]. This difficulty underscores the need to understand how specific kinds of visual and auditory artifacts shape users’ evaluations.
@@ 2.2 Heuristics and processing fluency in credibility judgments
Credibility refers to an individual’s judgment of the veracity of the content of communication [4]. Online environments complicate such judgments: information volume is high, source cues are sparse or ambiguous, and professional verification is limited [6, 28, 27, 38]. Under these conditions of overload and limited cognitive resources, people rely on heuristics; that is, readily accessible cues that can be evaluated quickly but are fallible [22, 15]. Social indicators such as endorsements, comments, or source reputation often serve this function [26], and popularity has been shown to be influential for credibility assessments explicitly in the context of deepfake content [17].
Deepfakes complicate credibility judgments further because they can appropriate trusted identities and obscure more diagnostic cues. Although synthetic videos may appear plausible, they often contain subtle irregularities. Such unexpected visual or auditory features create expectancy violations; deviations from what viewers assume recorded videos should look or sound like, which can reduce processing fluency, the subjective ease of processing a stimulus [48]. Processing fluency reliably influences credibility judgments across media: less legible text [35] and lower-quality audio [31] reduce perceived truthfulness [37].
People particularly rely on fluency when external knowledge is limited [3]. One proposed mechanism is familiarity: fluently processed information feels more familiar and is therefore judged as more likely to be true [8, 47]. In the context of deepfakes, this suggests that viewers may turn to the “feel” of ease or difficulty when evaluating unfamiliar videos or presenters. However, it remains unclear whether and how such fluency-based heuristics operate for videos, and especially for deepfake-generated faces and voices, where multiple perceptual channels may produce competing or ambiguous cues.
@@ 2.3 Credibility assessment as an interaction problem
In online environments, credibility judgments are shaped by an interactional process that includes the viewer and the design of digital interfaces. Users evaluate authenticity within feeds, video players, and recommendation systems that structure attention and impose cognitive constraints [15, 27]. HCI research has shown that interface-level cues such as modality, visual clarity, and system agency systematically influence heuristic credibility judgments [41, 39, 51]. In this perspective, deepfake artifacts can be understood as perceptual cues embedded in the interaction environment: they are surfaced to the user through the interface in the same way as other credibility-relevant signals. Studying how users respond to these artifacts therefore addresses the central question of how interactive systems mediate trust and uncertainty when synthetic media are encountered in real-world interface contexts.
@@ 4.4 Dependent variables
Following Graf et al. [12], subjective processing fluency was measured with a slider, with internally stored responses between 0 (difficult) and 100 (easy). This measure was used in all three studies.
@@
Based on the scale by Appelman and Sundar [4], message credibility was computed as the mean of three 7-point Likert items accurate, authentic, and believable, with response options ranging from 1 (“describes very poorly”) to 7 (“describes very well”). This scale was used in all three studies.
@@
In Study II and Study III, conceptual and factual learning were measured as the pre-to-post difference in a multiple choice questionnaire, following guidelines by Petersen et al. [34]. Pre- and post-stimulus knowledge questionnaires were administered with mixed item order and embedded attention checks.
@@ 4.5 Stimuli
An educational video was produced to serve as a common stimulus source for all three studies. The video was $447\text{\,}\mathrm{s}\mathrm{e}\mathrm{c}\mathrm{o}\mathrm{n}\mathrm{d}\mathrm{s}$ long and featured an actor presenting a script adapted from Petersen et al. [34]. All video editing was performed in Adobe After Effects 23.2.1 (see Appendix 2 and Appendix 3 for details on video processing). The original script consisted of sections on general virology, measles, Zika, and COVID-19; we excluded the COVID-19 section to avoid interference from prevalent, non-scientific interpretations of the disease.
@@
We used the Latent Image Animator with “vox-pt” weights by Wang and colleagues [45] to synthesize a deepfake version of the video. This approach requires a $256\times 256$ close-up video of a face and a $256\times 256$ portrait picture to generate a $256\times 256$ video of the portrait picture exhibiting the behavior captured in the video. The resulting animated portrait was then pasted into and aligned with the source video to obtain the deepfaked stimulus video. We synthesized a version of the presenter’s narration by fine-tuning a So-Vits-Svc-fork model [^2] https://github.com/voicepaw/so-vits-svc-fork. The model was trained on a $14.25\text{\,}\mathrm{m}\mathrm{i}\mathrm{n}\mathrm{u}\mathrm{t}\mathrm{e}\mathrm{s}$ long recording of a diverse set of English sentences adapted from the Harvard sentences [^3] https://harvardsentences.com. Once trained and provided with the audio of the source video, the model synthesized a version of the original narration in the new voice. The synthesized video stimulus was created by replacing the presenter’s face in the source video with the deepfake animation while retaining the original audio. The synthesized narration stimulus replaced the source video’s original narration track with the synthetic version. For the synthesized video and narration condition, we combined the deepfake stimulus video with the synthesized narration track.
@@ 5.1 Study I
More participants detected a digital alteration in the reduced visual clarity condition ($n=126$) than in the baseline condition ($n=22$). More participants were uncertain about digital alterations in the baseline group ($n=59$) than in the reduced visual clarity group ($n=32$). Similarly, more participants in the baseline group did not detect an alteration ($n=81$) compared to the reduced visual clarity group ($n=13$; see Figure 4). Building on work by Altay and Gilardi [2], who found that the belief of text being AI-generated negatively impacting its perceived credibility, we expanded on our preregistered analysis and examined whether believing that the video was digitally altered was associated with lower credibility. A Spearman rank correlation indicated a moderate negative association between alteration detection and credibility (r(331)=.25,p<.001r(331)=-.25,\,p<.001), see Appendix 7 for details.
@@ 5.3 Study III
Figure 7: Distribution of credible differences between the baseline video and the fully synthesized video. (a) shows the distribution of message credibility differences, (b) the distribution for processing fluency in Study III. Bayesian estimation [21] was used to calculate distributions of mean processing fluency and message credibility values for the baseline and synthesized video and narration conditions. The differences between the means are expressed in fractions of the pooled standard deviation (i.e., in units comparable to Cohen’s $d$). The 95% HDIs (High Density Intervals) indicate where the bulk of the credible differences fall. For processing fluency (b), the HDI is close to or includes 0, consistent with no meaningful difference between conditions. For message credibility (a), the HDI lies substantially above 0, and essentially all credible effect sizes exceed a small-effect threshold (Cohen’s $d=0.10$), indicating that the distance between the means is too large to accept $H_{0}$.
@@ 6 Discussion
In Study II, visual distortions and audio-video asynchrony reduced credibility compared to audio distortions alone but not the baseline, highlighting the role of sensory cues in credibility assessment. Study III found that only the combination of deepfake visuals and narration significantly reduced credibility compared to the baseline, suggesting a small additive effect. Contrary to Newman and Schwarz [31], we found no significant impact of audio manipulations on credibility.
@@
Groups of participants that indicated having and not having detected an alteration were present across all three studies, regardless of the presence of an manipulation. The data reveals a significant rate of inaccuracy in identifying distortions and deepfake videos, reflected in the high proportion of participants selecting “I don’t know.” These findings are consistent with Vaccari and Chadwick’s research, which links deepfakes to increased uncertainty [44]. Interestingly, the baseline condition in our studies highlights that the absence of manipulations does not reduce uncertainty. In Studies II and III, uncertainty was lowest in manipulation conditions and highest in baseline conditions without manipulations. Across studies, detecting (or believing to detect) an alteration was consistently associated with lower credibility. The negative correlations were small to moderate in Studies I and II and largest in Study III, indicating that perceived alteration reduced credibility regardless of whether manipulations were actually present.
@@
Our findings are in line with the association between source vividness and credibility as discussed by Lee and Shin in the context of deepfakes and misinformation [23]. We find, however, our source vividness measure to be more affected than our message credibility measure. This aligns with the understanding of video credibility assessment as complex and affected by multiple factors. While all participants learned from exposure to the stimulus, none of the employed manipulations affected their learning success. All learning-relevant information was mediated via the audio track of the videos. This suggests that the audio-based manipulations employed in Study II and Study III did not interfere with comprehensibility and that the measured reduction in source vividness and message credibility did not coincide with reduced apprehension.